OpenStudy (anonymous):

I'm trying to find critical numbers. The question is find any critical numbers of the function f(theta)=2 sec(theta)+tan(theta) 10 pi

OpenStudy (jango_in_dtown):

@Sophie_dofi

OpenStudy (jango_in_dtown):

\[2\sec \theta+\tan \theta \]

OpenStudy (jango_in_dtown):

??

OpenStudy (anonymous):

Yes

OpenStudy (jango_in_dtown):

so for critical numbers we have the derivative 0 or undefined right?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

I got I know it's something pi/6

OpenStudy (jango_in_dtown):

\[f'(\theta)=2\sec \theta \tan \theta+\sec ^{2}\theta\]

OpenStudy (jango_in_dtown):

\[=\sec \theta (2\tan \theta +\sec \theta)\]

OpenStudy (anonymous):

Yes I'm past that you end up with 2sin(x)=-1 and sin(x)=-1/2

OpenStudy (jango_in_dtown):

this expression is undefined or 0... so solve for sec theta=0 , cos theta=0 and 2tan theta + sec theta=0 or (1/2tan theta +sec theta)=0

OpenStudy (jango_in_dtown):

but sec theta never equals zero

OpenStudy (anonymous):

Is it 67pi/6

OpenStudy (anonymous):

71?

OpenStudy (jango_in_dtown):

cos theta =0 when theta is an odd multiple of pi/2

OpenStudy (anonymous):

I don't have a caululator... That's why I'm on here I left it at my moms

OpenStudy (hwyl):

wolframalpha.com

OpenStudy (anonymous):

Will not calculate equations like this.

OpenStudy (jango_in_dtown):

10pi < theta < 12 pi so here the solutions are given by

OpenStudy (jango_in_dtown):

\[10\pi<(2n+1)(\pi/2) <12\pi \]

OpenStudy (jango_in_dtown):

from here you will get two vaules of n, which are 10 and 11 .. so 21pi/2 and 23pi/2 are solution from cos theta=0

OpenStudy (jango_in_dtown):

next we need to compute the second part

OpenStudy (anonymous):

I'm taking notes with this bye I'm not ignoring you thank you very much

OpenStudy (anonymous):

*so

OpenStudy (jango_in_dtown):

and the second part gives us sin x=-1/2

OpenStudy (anonymous):

That's where I ended up

OpenStudy (jango_in_dtown):

in (10pi,12pi) we need the solution

OpenStudy (jango_in_dtown):

we know that sin 7pi/6 is -1/2

OpenStudy (anonymous):

10pi=31.42 and 12pi=37.69

OpenStudy (jango_in_dtown):

so we need to solve for \[10\pi<n \pi+ (-1)^{n}(7\pi/6)<12\pi\]

OpenStudy (anonymous):

And... Now I'm lost

OpenStudy (jango_in_dtown):

sin x=sin y implies \[x=n \pi+(-1)^{n}y\]

OpenStudy (jango_in_dtown):

I used this formula

OpenStudy (anonymous):

Ok

OpenStudy (jango_in_dtown):

see n cannot be greater than 12 or else the range goes out of 12pi and also n cannot be less than 10 or else the value will be less than 10pi

OpenStudy (jango_in_dtown):

check that n=10,11

OpenStudy (anonymous):

Oh

OpenStudy (jango_in_dtown):

so the solutions are 10pi+7pi/6 and 11pi-7pi/6

OpenStudy (jango_in_dtown):

that is 67/6pi and 59/6pi

OpenStudy (jango_in_dtown):

so the final solutions are 21pi/2, 23pi/2, 59pi/6 and 67pi/6

OpenStudy (anonymous):

lovely only 67pi/6 is an option

OpenStudy (anonymous):

Whohoo!

OpenStudy (jango_in_dtown):

ok then this is the solution.:)

OpenStudy (anonymous):

Thank you soooo much! Math is not my forte!

OpenStudy (jango_in_dtown):

do you understand the method???

OpenStudy (anonymous):

I have it all written down so I'll be able to try it again. I think I understand

OpenStudy (anonymous):

It's much easier with my caululator

OpenStudy (jango_in_dtown):

yes only if the calculator is allowed..:)

OpenStudy (anonymous):

My teacher allows it, we're a small class and were at the lowest level of calc available. She knows we won't ever go into a field with it or anything. As long as we understand the importance of learning ext.